# Intrinsic galaxy alignments in CAMELS simulations and the significant impact of baryon model

### v1

@@ Abstract
pubyear: 2025pagerange: Intrinsic galaxy alignments in CAMELS simulations and the significant impact of baryon modelB
@@ 1 Introduction
The intrinsic alignments of galaxies act both as a novel probe of galaxy formation and evolution (e.g. time_evolution_ia; Pk_TNG; sami_bh) and as a significant source of systematic error in weak gravitational lensing analyses (e.g. Hitara2004; krause_ia; amon; Secco_2022; KiDS25). Accurate modelling of intrinsic alignments is therefore essential – not only to extract cosmological information reliably from lensing surveys, but also to deepen our understanding of the processes affecting galaxy evolution.
@@
However, uncertainty in intrinsic alignment modelling remains a bottleneck for cosmological inference from state-of-the-art photometric galaxy surveys (Kirk_2010; Joachimi_2015; Paopiamsap_2024). It is, therefore, necessary to understand the mechanism of intrinsic alignments for different populations of galaxies under varying cosmological and astrophysical model assumptions. A further major source of uncertainty in weak lensing modelling are baryonic feedback processes, which also alter the matter distribution and impact observed signals. Measuring the relationship between intrinsic alignments and baryonic physics in simulations is a step towards mitigating modelling biases and, possibly, improving the containing power of cosmological datasets.
@@
Understanding the alignment mechanism is particularly important for realistic semi-empirical prescriptions, necessary for pipeline validation (lange_desi), field-level likelihood inference (borg), and simulation-based inference (2025A&A...694A.223V; sbi). But, even when using classical (less-constraining) techniques for cosmological inference, mismatches between assumed baryonic feedback models and intrinsic alignment prescriptions can introduce significant biases in cosmological parameter estimation, especially if they are not consistent with the properties of the particular galaxy sample.
There are well established theoretical frameworks for intrinsic alignments in different galaxy populations. For early-type or quiescent galaxies, observations (Singh_2015; 2019Johnston; 2021A&A...654A..76F) are broadly consistent with tidal-alignment models (catelan2001; Blazek_2015), where galaxy shapes align along the direction of the large-scale density field through tidal forces. In contrast, there is indeed observational evidence of galaxy spin axis alignments (Pen_2000; Mandelbaum2010; Jones2010; chisari_15). Recent studies indicate that these mechanisms can compete in high–star forming galaxies where baryonic feedback, mergers and morphology may shift the dominant alignment mechanisms (Joachimi_2015; Barsanti_2022; 10.1093/mnras/stad2728).
@@
In this work, we have both measured the intrinsic shapes of galaxies across 1000 CAMELS simulations (Villaescusa_Navarro_2021) and detected their alignment with large-scale structure. By cross-correlating the projected galaxy ellipticities with the underlying dark matter, we sidestep considerations of galaxy bias to obtain a clean two-point alignment signal.
@@
Previous analyses have measured intrinsic alignments in simulations, but with fixed (or restricted) cosmological and astrophysical simulation parameter values (e.g 2015SSRv..193...67K; chisari_15; 2020Zjupa; 2021MNRAS.501.5859T; samuroff_sim; Delgado_2023).
@@
Figure 1: A $1.67\,h^{-1}\,\text{Mpc}$ thick slice of the CAMELS LH-310 simulation viewed along the $y$-axis. The background shows the log-normal dark matter density with stellar density overlaid. Galaxy shapes are represented as ellipses with eccentricity and orientation encoding the magnitude and angle of the measured ellipticity $\varepsilon$. On average, galaxy shapes tend to align preferentially with the principal axes of the dark matter distribution.
@@ 2 Intrinsic alignment theory
We follow the approach taken by direct detection measurements from galaxy survey data (Mandelbaum2006; 2011A&A...527A..26J; 2019Johnston; 2021A&A...654A..76F; 2023MNRAS.524.2195S; siegel2025) by measuring alignments via the projected cross-correlation function $w_{m+}$ as a function of the comoving transverse separation $r_{\perp}$. Note that we use the subscript $m$ notation to refer to our correlation with respect to matter particles in the simulation, rather than galaxies $g$. This is given as a radial projection of the anisotropic two-point correlation function:
@@
for perpendicular $k_{\perp}$ and $k_{z}$ radial Fourier components. This assumes no redshift-space distortions (1987MNRAS.227....1K) or shape projection effects (see 2015MNRAS.450.2195S).
In this work, we do not explicitly model $P_{\delta I}$ – the aim is provide a model-agnostic detection of the signal and its dependence on the cosmological and astrophysical parameters. However, in the literature there are a number of popular proposed prescriptions for $P_{\delta I}$, including the linear and non-linear alignment (NLA) models (catelan2001; Hitara2004; Bridle_2007), the tidal alignment and tidal torquing (TATT) model (2019PhRvD.100j3506B), and halo-based models (2021MNRAS.501.2983F).
@@ 3.1 The CAMELS simulations
The CAMELS simulations [^1] https://www.camel-simulations.org/ are sets of hydrodynamical and N-body cosmological simulations that sample a high-dimensional parameter volume in order to explore connections between cosmology and astrophysical processes. We measure galaxy shapes in the Latin-Hypercube (LH) set of simulations that use the hydrodynamic IllustrisTNG code. This set contains 1000 simulations, each having different values of selected cosmological and astrophysical parameters as detailed in Table 1. The IllustrisTNG cosmic structure and galaxy formation model uses the AREPO gravity code (springel2010; Vogelsberger2012) and incorporates various astrophysical processes such as baryonic effects detailed in Vogelsberger2013; Pillepich2017.
All hydrodynamical simulations in CAMELS contain $256^{3}$ dark matter particles and $256^{3}$ fluid elements, and follow time evolution from redshift $z=127$ down to $z=0$ in a periodic box of volume $(25\,h^{-1}\,\text{Mpc})^{3}$. The simulations also assume a spatially flat cold dark matter (CDM) cosmological model with constant parameters: baryon density $\Omega_{b}=0.049$, Hubble parameter $h=0.6711$, dark matter equation of state w=1w=-1, and sum of neutrino masses $\sum m_{\nu}=0$. While the zero neutrino mass is unphysical, we expect this common approximation to have negligible impact on our results. Further details on the effects of these parameters on the simulations can be found in Villaescusa-Navarro_2023.
@@ 3.2 Simulated shape catalogue
For each simulation, stellar and dark matter particle positions at redshift $z=0$ were recorded together with the subhalo catalogue identified by the SUBFIND algorithm (Dolag_Springer; springel_2001). Galaxies were selected from subhaloes based on the following criterion:
@@
Here $Q_{pq}$ is defined similarly to Eq. 4.2 from BARTELMANN2001291, but using stellar density instead of brightness (flux):
@@
Dark matter and galaxy positions from our catalogues were transformed from Cartesian coordinates to right ascension and declination (RA, DEC) coordinates by positioning the centre of the simulation box at $10,000\,h^{-1}$ Mpc co-moving distance from the simulated observer. This setup enables the use of the two-point correlation function measurement pipeline developed in 2019Johnston, which employs TreeCorr (treecorr) to compute the projected correlation functions $w_{m+}(r_{\perp})$, $w_{m\times}(r_{\perp})$, and $w_{mm}(r_{\perp})$. The coordinate transformation was performed to implement the pipeline and the positioning value of $10,000\,h^{-1}$ Mpc was chosen to minimise the error between separations in RA DEC coordinates and the Cartesian projected separations [^2] The coordinate transform generates a $0.3\%$ maximum distance error relative to box size..
@@ 3.3 Correlation functions
(Peebles1974), where $DD$ is the number of dark matter–dark matter pairs in bins of transverse and line-of-sight separation $(r_{\perp},\Pi)$, and $RR$ is the number of such pairs in a corresponding randomly generated catalogue. Random catalogues are constructed with the same number of simulated dark matter particles as the original data. Given this high number of particles, the Poisson noise is much smaller than the signal. We do not use the Landy-Szalay estimator as our catalogues are spatially uniform and the depth of variation becomes very small for the simulation box placed at high redshift.
We compute the correlation between dark matter and intrinsic shear using an estimator similar to that in Mandelbaum2006:
@@
We considered pairwise signals for the 2D correlation functions (Eqs. 78 and 10) in co-moving range $|\Pi|\leq 12\,h^{-1}$ Mpc discretized into 24 bins of $\Delta\Pi=1\,h^{-1}$ Mpc. We note that $\Pi_{\text{max}}$ is much smaller than typical $\Pi_{\text{max}}$ values used in surveys due to our limited simulation volume (Mandelbaum2006; 2011A&A...527A..26J; 2019Johnston). Transverse separations $r_{\perp}$ were considered in six logarithmically spaced bins in the range $0.3\leq r_{\perp}\leq 20\,h^{-1}$ Mpc. We sum over $\Delta\Pi$ intervals to approximate the integral in Eq. 1 and obtain projected correlation functions $w_{m+}(r_{\perp})$, $w_{m\times}(r_{\perp})$ and $w_{mm}(r_{\perp})$.
With our estimate of $w_{mm}$, the small volume of the CAMELS simulations induces an additive bias through the integral constraint (Peebles:1980yev) – this is a common observation effect in small galaxy surveys (e.g. 1993MNRAS.263..360R; 1999MNRAS.306..988M). In our case, if uncorrected, we measure negative values for $w_{mm}$ at high $r_{\perp}$. Our correction only affects results that use the ratio $\frac{w_{m+}}{w_{mm}}$, and we find that the conclusions are not affected by the exact value.
@@
We split the galaxy catalogues at a specific star formation rate (sSFR) of $10^{-10.5}$ /yr (L_Bisigello_Euclid_split) and make three samples: full, star forming, and quiescent. Note that in IllustrisTNG, the galactic SFR is defined as the sum of star formation rates of all gas cells existing in the subhalo.
@@ 4.1 Alignment signal from catalogues
We also demonstrate that the alignments of quiescent galaxies are significantly stronger than those of high star forming galaxies in our CAMELS catalogues. Again in Figure 2, $w_{m+}(r_{\perp})$ for the sample of galaxies below sSFR $10^{-10.5}$ yr-1 is consistently higher than that of the high star forming sample. This agrees with previous studies showing that alignments behave differently in early and late-type galaxies (hirata2007; mandelbaum2011; samuroff_sim). We note that the star-forming population may be contaminated by late-type galaxies as some models of IA (chisari_15; 2019PhRvD.100j3506B) predict a negative (tangential) alignment for high star-forming galaxies.
Figure 3 presents mean measured orientation-only correlation functions, denoted $\tilde{w}_{m+}(r_{\perp})$ calculated from orientation-only catalogues outlined in Section 3.4. As before, mean functions from the full, quiescent and star forming samples are taken. We find that normalizing ellipticity magnitudes affects alignments uniformly across $0.3\leq r_{\perp}\leq 20\,h^{-1}$ Mpc as the functions $w_{m+}(r_{\perp})$ and $\tilde{w}_{m+}(r_{\perp})$ (in Figures 2 and 3 respectively) have similar scale dependence. This is a feature found in intrinsic alignment models (Schneider_2010; Blazek_2011) where the scale dependence is set by the tidal field. This result has also been observed previous works on surveys such as 2024MNRAS.534..3540.
@@ 4.2 Validation
In line with previous studies of intrinsic alignments (Mandelbaum2006; hirata2007; 2011A&A...527A..26J; 2019Johnston), we validate that the detected alignment signal arises from astrophysical effects rather than systematic errors introduced by the measurement pipeline or simulation setup.
We compute the mean position shape cross-component correlation function $w_{m\times}(r_{\perp})$, alongside $w_{m+}(r_{\perp})$, as a null test. This statistic corresponds to rotating all catalogue galaxies’ shapes by 45 degrees and recomputing $w_{m+}(r_{\perp})$. Since $w_{m\times}(r_{\perp})$ changes sign under parity (rotating galaxies clockwise or anti-clockwise), it must be zero to preserve the parity invariance of galaxy formation. The measured $w_{m\times}(r_{\perp})$ for both full and split galaxy samples is shown in Figure 4, with error bars denoting the standard error on the mean. All measurements are consistent with zero, indicating no significant systematic bias in our $w_{m+}$ signal. We note the presence of B-modes may provide a small, non-zero $w_{m\times}(r_{\perp})$ signals, which has been proposed as possibility for intrinsic alignments (Georgiou_2024), which we do not detect in these simulation.

### v2

@@ Abstract
pubyear: 2026pagerange: Intrinsic galaxy alignments in CAMELS simulations and the significant impact of baryon modelB
@@ 1 Introduction
The intrinsic alignments of galaxies act both as a novel probe of galaxy formation and evolution (e.g. Schmitz et al., 2018; Shi et al., 2021; Barsanti et al., 2023b) and as a significant source of systematic error in weak gravitational lensing analyses (e.g. Hirata and Seljak, 2004; Krause et al., 2016; Amon et al., 2022; Secco et al., 2022; Wright et al., 2025). Accurate modelling of intrinsic alignments is therefore essential – not only to extract cosmological information reliably from lensing surveys, but also to deepen our understanding of the processes affecting galaxy evolution.
@@
However, uncertainty in intrinsic alignment modelling remains a bottleneck for cosmological inference from state-of-the-art photometric galaxy surveys (Kirk et al., 2010; Joachimi et al., 2015; Paopiamsap et al., 2024). It is, therefore, necessary to understand the mechanism of intrinsic alignments for different populations of galaxies under varying cosmological and astrophysical model assumptions. A further major source of uncertainty in weak lensing modelling are baryonic feedback processes, which also alter the matter distribution and impact observed signals. Measuring the relationship between intrinsic alignments and baryonic physics in simulations is a step towards mitigating modelling biases and, possibly, improving the containing power of cosmological datasets.
@@
Understanding the alignment mechanism is particularly important for realistic semi-empirical prescriptions, necessary for pipeline validation (Lange et al., 2024), field-level likelihood inference (Porqueres et al., 2023), and simulation-based inference (von Wietersheim-Kramsta et al., 2025; Jeffrey et al., 2025). But, even when using classical (less-constraining) techniques for cosmological inference, mismatches between assumed baryonic feedback models and intrinsic alignment prescriptions can introduce significant biases in cosmological parameter estimation, especially if they are not consistent with the properties of the particular galaxy sample.
There are well established theoretical frameworks describing intrinsic alignments in different galaxy populations. For early-type, quiescent galaxies, observations demonstrate significant correlations between galaxy shapes and the density field (Singh et al., 2015b; Johnston et al., 2019; Fortuna et al., 2021b), broadly consistent with tidal-alignment models where stellar distributions are elongated along the direction of surrounding overdensities (Catelan et al., 2001; Hirata and Seljak, 2004; Blazek et al., 2015). In contrast, late-type or star-forming galaxies are expected to align primarily through angular momentum acquisition: tidal torques during halo formation generate galaxy spin directions that are tangentially correlated with the surrounding tidal field (Doroshkevich, 1970; White, 1984). Observational studies have found evidence for correlated spin axes in spiral galaxies (Pen et al., 2000; Mandelbaum et al., 2010a; Jones et al., 2010; Chisari et al., 2015), although corresponding two-point correlations of projected galaxy shapes are significantly weaker (Mandelbaum et al., 2010b; Johnston et al., 2019), reflecting the lower amplitude of tidal-torque-induced intrinsic alignments. Motivated by early observations, Crittenden et al. (2001) introduced a quadratic alignment model for spiral galaxies, modelling systems as thin discs whose observed ellipticities arise from angular-momentum-induced couplings to the large-scale tidal field. While Blazek et al. (2019b) proposed a perturbative approach that includes both the tidal shearing and tidal torquing alignment mechanism. Recent work (Joachimi et al., 2015; Barsanti et al., 2022, 2023a) suggests that baryonic feedback, mergers, and morphological evolution may blur the distinction between alignment mechanisms in early- and late-type galaxies.
@@
In this work, we have both measured the intrinsic shapes of galaxies across 1000 CAMELS simulations (Villaescusa-Navarro et al., 2021) and detected their alignment with large-scale structure. By cross-correlating the projected galaxy ellipticities with the underlying dark matter, we sidestep considerations of galaxy bias to obtain a clean two-point alignment signal.
@@
Previous analyses have measured intrinsic alignments in simulations, but with fixed (or restricted) cosmological and astrophysical simulation parameter values (e.g Kiessling et al., 2015; Chisari et al., 2015; Zjupa et al., 2020; Tenneti et al., 2021; Samuroff et al., 2021; Delgado et al., 2023).
@@
Figure 1: A $1.67\,h^{-1}\,\text{Mpc}$ thick slice of the CAMELS LH-310 simulation viewed along the $y$-axis. The background shows the log-normal dark matter density with stellar density overlaid. Galaxy shapes are represented as ellipses with eccentricity and orientation encoding the magnitude and angle of the measured ellipticity $\varepsilon$. On average, galaxy shapes tend to align preferentially with the principal axes of the dark matter distribution, though intrinsic alignments in observational data are only detected statistically and not at the level of individual galaxies.
@@ 2 Intrinsic alignment theory
We follow the approach taken by direct detection measurements from galaxy survey data (Mandelbaum et al., 2006; Joachimi et al., 2011; Johnston et al., 2019; Fortuna et al., 2021b; Samuroff et al., 2023; Siegel et al., 2025) by measuring alignments via the projected cross-correlation function $w_{m+}$ as a function of the comoving transverse separation $r_{\perp}$. Note that we use the subscript $m$ notation to refer to our correlation with respect to matter particles in the simulation, rather than galaxies $g$. This is given as a radial projection of the anisotropic two-point correlation function:
@@
for perpendicular $k_{\perp}$ and $k_{z}$ radial Fourier components. This assumes no redshift-space distortions (Kaiser, 1987) or shape projection effects (see Singh et al., 2015a).
In this work, we do not explicitly model $P_{\delta I}$ – the aim is provide a model-agnostic detection of the signal and its dependence on the cosmological and astrophysical parameters. However, in the literature there are a number of popular proposed prescriptions for $P_{\delta I}$, including the linear and non-linear alignment (NLA) models (Catelan et al., 2001; Hirata and Seljak, 2004; Bridle and King, 2007), the tidal alignment and tidal torquing (TATT) model (Blazek et al., 2019a), and halo-based models (Fortuna et al., 2021a).
@@ 3.1 The CAMELS simulations
The CAMELS simulations [^1] https://www.camel-simulations.org/ are sets of hydrodynamical and N-body cosmological simulations that sample a high-dimensional parameter volume in order to explore connections between cosmology and astrophysical processes. We measure galaxy shapes in the Latin-Hypercube (LH) set of simulations that use the hydrodynamic IllustrisTNG code. This set contains 1000 simulations, each having different values of selected cosmological and astrophysical parameters as detailed in Table 1. The IllustrisTNG cosmic structure and galaxy formation model uses the AREPO gravity code (Springel, 2010; Vogelsberger et al., 2012) and incorporates various astrophysical processes such as baryonic effects detailed in Vogelsberger et al. (2013); Pillepich et al. (2017).
All hydrodynamical simulations in CAMELS contain $256^{3}$ dark matter particles and $256^{3}$ fluid elements, and follow time evolution from redshift $z=127$ down to $z=0$ in a periodic box of volume $(25\,h^{-1}\,\text{Mpc})^{3}$. The simulations also assume a spatially flat cold dark matter (CDM) cosmological model with constant parameters: baryon density $\Omega_{b}=0.049$, Hubble parameter $h=0.6711$, dark matter equation of state w=1w=-1, and sum of neutrino masses $\sum m_{\nu}=0$. While the zero neutrino mass is unphysical, we expect this common approximation to have negligible impact on our results. Further details on the effects of these parameters on the simulations can be found in Villaescusa-Navarro et al. (2023).
@@ 3.2 Simulated shape catalogue
For each simulation, stellar and dark matter particle positions at redshift $z=0$ were recorded together with the subhalo catalogue identified by the SUBFIND algorithm (Dolag et al., 2009; Springel et al., 2001). Galaxies were selected from subhaloes based on the following criterion:
@@
Here $Q_{pq}$ is defined similarly to Eq. 4.2 from Bartelmann and Schneider (2001), but using stellar density instead of brightness (flux):
@@
Dark matter and galaxy positions from our catalogues were transformed from Cartesian coordinates to right ascension and declination (RA, DEC) coordinates by positioning the centre of the simulation box at $10,000\,h^{-1}$ Mpc co-moving distance from the simulated observer. This setup enables the use of the two-point correlation function measurement pipeline developed in Johnston et al. (2019), which employs TreeCorr (Jarvis, 2015) to compute the projected correlation functions $w_{m+}(r_{\perp})$, $w_{m\times}(r_{\perp})$, and $w_{mm}(r_{\perp})$. The coordinate transformation was performed to implement the pipeline and the positioning value of $10,000\,h^{-1}$ Mpc was chosen to minimise the error between separations in RA DEC coordinates and the Cartesian projected separations [^2] The coordinate transform generates a $0.3\%$ maximum distance error relative to box size..
@@ 3.3 Correlation functions
(Peebles and Hauser, 1974), where $DD$ is the number of dark matter–dark matter pairs in bins of transverse and line-of-sight separation $(r_{\perp},\Pi)$, and $RR$ is the number of such pairs in a corresponding randomly generated catalogue. Random catalogues are constructed with the same number of simulated dark matter particles as the original data. Given this high number of particles, the Poisson noise is much smaller than the signal. We do not use the Landy-Szalay estimator as our catalogues are spatially uniform and the depth of variation becomes very small for the simulation box placed at high redshift.
We compute the correlation between dark matter and intrinsic shear using an estimator similar to that in Mandelbaum et al. (2006):
@@
We considered pairwise signals for the 2D correlation functions (Eqs. 78 and 10) in co-moving range $|\Pi|\leq 12\,h^{-1}$ Mpc discretized into 24 bins of $\Delta\Pi=1\,h^{-1}$ Mpc. We note that $\Pi_{\text{max}}$ is much smaller than typical $\Pi_{\text{max}}$ values used in surveys due to our limited simulation volume (Mandelbaum et al., 2006; Joachimi et al., 2011; Johnston et al., 2019). Transverse separations $r_{\perp}$ were considered in six logarithmically spaced bins in the range $0.3\leq r_{\perp}\leq 20\,h^{-1}$ Mpc. We sum over $\Delta\Pi$ intervals to approximate the integral in Eq. 1 and obtain projected correlation functions $w_{m+}(r_{\perp})$, $w_{m\times}(r_{\perp})$ and $w_{mm}(r_{\perp})$.
With our estimate of $w_{mm}$, the small volume of the CAMELS simulations induces an additive bias through the integral constraint (Peebles, 1980) – this is a common observation effect in small galaxy surveys (e.g. Roche et al., 1993; Magliocchetti and Maddox, 1999). In our case, if uncorrected, we measure negative values for $w_{mm}$ at high $r_{\perp}$. Our correction only affects results that use the ratio $\frac{w_{m+}}{w_{mm}}$, and we find that the conclusions are not affected by the exact value.
@@
We split the galaxy catalogues at a specific star formation rate (sSFR) of $10^{-10.5}$ /yr (Bisigello et al., 2020) and make three samples: full, star forming, and quiescent. Note that in IllustrisTNG, the galactic SFR is defined as the sum of star formation rates of all gas cells existing in the subhalo.
@@ 4.1 Alignment signal from catalogues
We also demonstrate that the alignments of quiescent galaxies are significantly stronger than those of high star forming galaxies in our CAMELS catalogues. Again in Figure 2, $w_{m+}(r_{\perp})$ for the sample of galaxies below sSFR $10^{-10.5}$ yr-1 is consistently higher than that of the high star forming sample. This agrees with previous studies showing that alignments behave differently in early and late-type galaxies (Hirata et al., 2007; Mandelbaum et al., 2010b; Samuroff et al., 2021). We note that the star-forming population may be contaminated by late-type galaxies as some models of IA (Chisari et al., 2015; Blazek et al., 2019a) predict a negative (tangential) alignment for high star-forming galaxies.
Figure 3 presents mean measured orientation-only correlation functions, denoted $\tilde{w}_{m+}(r_{\perp})$ calculated from orientation-only catalogues outlined in Section 3.4. As before, mean functions from the full, quiescent and star forming samples are taken. We find that normalizing ellipticity magnitudes affects alignments uniformly across $0.3\leq r_{\perp}\leq 20\,h^{-1}$ Mpc as the functions $w_{m+}(r_{\perp})$ and $\tilde{w}_{m+}(r_{\perp})$ (in Figures 2 and 3 respectively) have similar scale dependence. This is a feature found in intrinsic alignment models (Schneider and Bridle, 2010; Blazek et al., 2011) where the scale dependence is set by the tidal field. This result has also been observed previous works on surveys such as Lamman et al. (2024).
@@ 4.2 Validation
In line with previous studies of intrinsic alignments (Mandelbaum et al., 2006; Hirata et al., 2007; Joachimi et al., 2011; Johnston et al., 2019), we validate that the detected alignment signal arises from astrophysical effects rather than systematic errors introduced by the measurement pipeline or simulation setup.
We compute the mean position shape cross-component correlation function $w_{m\times}(r_{\perp})$, alongside $w_{m+}(r_{\perp})$, as a null test. This statistic corresponds to rotating all catalogue galaxies’ shapes by 45 degrees and recomputing $w_{m+}(r_{\perp})$. Since $w_{m\times}(r_{\perp})$ changes sign under parity (rotating galaxies clockwise or anti-clockwise), it must be zero to preserve the parity invariance of galaxy formation. The measured $w_{m\times}(r_{\perp})$ for both full and split galaxy samples is shown in Figure 4, with error bars denoting the standard error on the mean. All measurements are consistent with zero, indicating no significant systematic bias in our $w_{m+}$ signal. We note the presence of B-modes may provide a small, non-zero $w_{m\times}(r_{\perp})$ signals. It has been proposed that intrinsic alignments may be sensitive to such B-mode contributions (Georgiou et al., 2024), which we do not detect in these simulations.